THE B(1/2^{2}P_{3/2}) → X(1/2^{2}Σ^{+}) TRANSITION IN XeBr

\underline{JASON O. CLEVENGER}, Department of Chemistry, MIT, Cambridge, MA 02139; and JOEL TELLINGHUISEN, Department of Chemistry, Vanderbilt University, Nashville, Tennessee 37235.

The B(1/2^{2}P_{3/2}) → X(1/2^{2}Σ^{+}) transition in XeBr was recorded in high resolution, using a CCD array detector to record spectra from Tesla discharge sources containing isotopically pure ^{136}Xe with ^{81}Br_{2} or ^{79}Br_{2}^{a}^{b}. The high signal/noise capabilities of the detector permitted the measurement of discrete vibrational structure in this system, which has normally been treated as a purely bound-free transition. The assignments comprised 119 v′-v′′ bands for ^{136}Xe^{81}Br and 86 for ^{136}Xe^{79}Br, spanning v′= 0-33 and v′′= 0-16. The van der Waals ground state was analyzed through fits to the customary polynomials and to near- dissociation expansions. Franck-Condon calculations were used to locate the X-state potential on the internuclear axis relative to the B state, which was modeled as a Rittner potential. The following fundamental spectroscopic constants (units cm^{−1}, for ^{136}Xe^{81}Br) were obtained from the analysis: T′_{e} = 35 863.2, ω′_{e} = 135.72, ω_{e}x′_{e} = 0.32, ω′′_{e} = 25.7, ω_{e}x′_{e} = 0.62. The ground state has a dissociation energy D′′_{e} = 254 ± 2 cm^{−1}and supports 24 bound vibrational levels.

aJason O. Clevenger and Joel Tellinghuisen Chem. Phys. Lett. \underline{231}, 515 December 1994.

bJason O. Clevenger and Joel Tellinghuisen J. Chem. Phys. \underline{103}(22), 9611 December 1995.